OpenStudy (anonymous):

sqrt x - 98 + 8 = 20 Does this have an extraneous solution? Steps for solving please!

OpenStudy (anonymous):

Looks like this : \[\sqrt{x - 98} + 8 = 20\]

OpenStudy (anonymous):

@Yttrium

OpenStudy (anonymous):

@Jack1 @austinL

OpenStudy (yttrium):

Can you define extraneous solution?

OpenStudy (anonymous):

@Yttrium "A solution of a simplified version of an equation that does not satisfy the original equation." - http://www.mathwords.com/e/extraneous_solution.htm

OpenStudy (yttrium):

Okay. Now you know it's meaning. Can you tell me what to do first?

OpenStudy (anonymous):

Honestly, I was poorly introduced to square roots so I was never taught the steps, though I'd assume you work outside the square root first?

OpenStudy (yttrium):

Okay. What you actually need to do is to isolate the terms having square roots. So, what will happen in our given equation?

OpenStudy (anonymous):

Subtract 8. We'll have \[\sqrt{x - 98} = 12\]

OpenStudy (yttrium):

Correct. So what will happen next?

OpenStudy (anonymous):

I assume we eliminate the 98 which would become this \[\sqrt{x} = 110\]

OpenStudy (yttrium):

Ooopss. You can not easily eliminate 98 there. You must square first the whole equation to cancel the sqrt sign and then do the transpose.

OpenStudy (anonymous):

And how would I do that?

OpenStudy (yttrium):

We have \[\sqrt{x-98} = 12]\] right? So, \[[\sqrt{x-98} = 12]^2\] And this will give us x-98 = 144

OpenStudy (anonymous):

How does that work?! I don't even see how you squared the whole equation like that. But as for what you showed me, we'd then work with the 98 and make it this : \[x = 242\]

OpenStudy (yttrium):

It's just like \[[\sqrt{x-98}]=(12)^2\] Yes. So, that will be the answer. And through substitution, you can chekc whether it is extraneous or not. :)

OpenStudy (anonymous):

Hmm.. now I have that in my notes. As for the equation, it wouldn't be extraneous since we have a solution, correct?

OpenStudy (anonymous):

@Yttrium

OpenStudy (yttrium):

Yes! :)

OpenStudy (anonymous):

Alright thanks!

OpenStudy (yttrium):

No prob. :))

OpenStudy (yttrium):

Hope, you understand now. It's a pleasure for me. :))